# Robust $Q$-Learning for finite ambiguity sets

### v1

@@ 1.3. Structure
The remainder of this paper is as follows. In Section 2 we present the setting and the associated stochastic control problem. A numerical solution to this optimization problem via $Q$-learning is presented in Section 3 together with our main result guaranteeing convergence of the $Q$-learning algorithm. In Section 4 we provide numerical experiments showcasing the tractability of our $Q$-learning algorithm. The proofs are reported in Section 5.
@@ 2. Setting and specification of the problem
In this section we introduce the underlying framework which we use to establish a robust Q-learning algorithm for finite ambiguity sets.
@@ 2.1. Setting
where for all $k\in\{1,\dots,N\}$ and for all $(x,a)\in\mathcal{X}\times A$ we have that $\mathbb{P}^{(k)}(x,a)\in\mathcal{M}_{1}(\mathcal{X})$. Note that in the degenerate case where $\mathcal{P}(x,a)$ contains only a single probability distribution, we get a classical Markov decision process (i.e. a non-robust one), compare, e.g., [7].
@@
where the notation $\mathbb{P}=\delta_{x}\otimes\mathbb{P}_{0}\otimes\mathbb{P}_{1}\otimes...\in\mathfrak{P}_{x,a}$ abbreviates
@@
The construction of ambiguity sets $\mathfrak{P}_{x,\textbf{a}}$ provided in (2.1) and (2.2) is in the related literature often referred to as a $(s,a)$-rectangular ambiguity set, (compare [17] and [48]) and enables to obtain a dynamic programming principle of the associated Markov decision problem.
After execution of an action $a_{t}(X_{t})$, the agent receives a feedback on the quality of the chosen action in terms of a reward $r(X_{t},a_{t}(X_{t}),X_{t+1})$ where $r$ denotes some reward function $r:\mathcal{X}\times A\times\mathcal{X}\rightarrow\mathbb{R}$.
@@ 2.3. Specification of the optimization problem
The robust optimization problem consists, for every initial state $x\in\mathcal{X}$, in maximizing the expected value of $\sum_{t=0}^{\infty}\alpha^{t}r(X_{t},a_{t}(X_{t}),X_{t+1})$ under the worst case measure from $\mathfrak{P}_{x,\textbf{a}}$ over all possible policies $\textbf{a}\in\mathcal{A}$, where $\alpha\in\mathbb{R}$ is some discount factor accounting for time preferences of rewards. Therefore, the value function
@@ 2.4. Dynamic Programming
We consider the one step optimization problem
@@
(2.5) $\mathcal{X}\times A\ni(x,a)\mapsto Q^{*}(x,a):=\inf_{\mathbb{P}\in\mathcal{P}(x,a)}\mathbb{E}_{\mathbb{P}}\left[r(x,a,X_{1})+\alpha V(X_{1})\right].$
We can interpret $Q^{*}(x,a)$ as the quality of executing action $a$ when in state $x$ (under the worst case measure) with the best possible action leading to $V(x)$. By using the above definitions we indeed obtain the following dynamic programming equation, also famously known in the non-robust formulation as Bellman equation, compare e.g. [12, p.164] for a non-robust version.
@@ 3. Robust Q-Learning Algorithm
4. Examples
@@ 4.1. Coin Toss
As in [28, Example 4.1] we consider an agent playing a coin toss game.
@@
At each time step $t$ the agent can make a bet whether the sum of the next throw strictly exceeds the previous sum (i.e. $X_{t+1}>X_{t}$), or whether it is strictly smaller (i.e. $X_{t+1}<X_{t}$). If the agent is correct, she gets $1\$$, however if the agent is wrong she has to pay $1\$$. The agent also has the possibility not to play. We model this by considering the following reward function:
@@
which are Wasserstein-balls around the reference probability $\operatorname*{Bin}(10,0.5)$ with radii $1$ and $2$, respectively. As one can see, e.g., from [27, Equation (4.2)], $\mathcal{P}^{3}$ is the smallest Wasserstein-ball around the reference measure that contains $\operatorname*{Bin}(10,0.6)$, whereas $\mathcal{P}^{4}$ is the smallest Wasserstein-ball containing $\operatorname*{Bin}(10,0.3)$.
@@ 4.2. Stock investing
This example is a variant of [28, Example 4.3] showcasing the use of tailored ambiguity sets which are not possible to construct and use in the same manner with Wasserstein or Entropic ambiguity sets as they are considered in [6], [22], and [44].
@@ 5.1. Auxiliary Results
Let $10<\alpha<$ 1, and let the ambiguity set $\mathcal{P}$ be defined in (2.1). Then the following fixed-point equation holds true for the optimal $Q$-value function defined in (2.5):
@@
This follows directly by definition of $Q^{*}$ and by Proposition 2.1, compare also [27, Proof of Lemma 17]. ∎
@@
The proof follows is analogue to [27, Proof of Lemma 18]. ∎
The following lemma is a result from stochastic approximation theory adjusted to our setting, compare [27, Lemma A.1] for the presented formulation adjusted to the underlying setting. We refer also to [36, Lemma 1], [14] [18, Theorem 1], [38, Lemma 12], and [42, Lemma 3] for further reference.
@@
5.2. Proofs of the results from Section 2 and Section 3
@@
Note that the equality $\max_{a\in A}Q^{*}(x,a)=\mathcal{T}V(x)$ directly follows by definition of the operator $\mathcal{T}$, and it only remains to show $\mathcal{T}V=V$. To this end, we aim at applying [28, Theorem 3.1] with $p=0$ and verify its assumptions, i.e., we verify [28, Assumption 2.2] and [28, Assumption 2.4] by showing the following four properties.
@@
To see that (i) holds, we note that $\mathcal{P}$ is by definition non-empty and since the set $\mathcal{P}(x,a)$ is finite it is also compact for all $(x,a)\in\mathcal{X}\times A$. To show the upper hemicontinuity, we consider some $(x,a)\in\mathcal{X}\times A$ and we let $(x_{n},a_{n},\mathbb{P}_{n})_{n\in\mathbb{N}}\subseteq\operatorname{Gr}(\mathcal{P}):=\{(x,a,\mathbb{P})~{}|~{}(x,a)\in\mathcal{X}\times A,\mathbb{P}\in\mathcal{P}(x,a)\}$ be a sequence such that $(x_{n},a_{n})\rightarrow(x,a)\in\mathcal{X}\times A$ as $n\rightarrow\infty$. Since both, $\mathcal{X}$ and $A$ are finite, we have $(x_{n},a_{n})=(x,a)$ for $n$ large enough. Hence, $\mathbb{P}_{n}\in\mathcal{P}(x,a)$ for $n$ large enough. As $\mathcal{P}(x,a)$ is finite we also have that $\mathbb{P}_{n}=\mathbb{P}^{(i)}(x,a)$ for infinitely many indices $n$ for some $i\in\{1,\dots,N\}$. Hence, we can find a subsequence $(\mathbb{P}_{n_{k}})_{k\in\mathbb{N}}$ such that $\mathbb{P}_{n_{k}}\rightarrow\mathbb{P}^{(i)}(x,a)\in\mathcal{P}(x,a)$ weakly as $k\rightarrow\infty$ and the upper hemicontinuity of $\mathcal{P}$ follows by the characterization provided in [2, Theorem 17.20].
To show the lower hemicontinuity let $(x,a)\in\mathcal{X}\times A$, $\mathbb{P}\in\mathcal{P}(x,a)$ and consider a sequence $(x_{n},a_{n})_{n\in\mathbb{N}}\subseteq\mathcal{X}\times A$ with $(x_{n},a_{n})\rightarrow(x,a)$ as $n\rightarrow\infty$. By definition of $\mathcal{P}$, there exists some $i\in\{1,\dots,N\}$ such that $\mathbb{P}^{(i)}(x,a)=\mathbb{P}$. We use this observation to define
$\mathbb{P}_{n}:=\mathbb{P}^{(i)}(x_{n},a_{n})\in\mathcal{P}(x_{n},a_{n}),~{}n\in\mathbb{N},$
@@
$k_{t}^{*}\in\operatorname*{arg\,min}_{k=1,...,N}\mathbb{E}_{\mathbb{P}^{(k)}(X_{t},a_{t}(X_{t}))}\left[r(X_{t},a_{t}(X_{t}),X_{t+1})+\alpha\max_{b\in A}Q_{t}(X_{t+1},b)\right]$

### v2

@@ 1.3. Structure
The remainder of this paper is as follows. In Section 2 we present the setting and the associated stochastic control problem. A numerical solution to this optimization problem via $Q$-learning is presented in Section 3 together with our main result guaranteeing convergence of the $Q$-learning algorithm. Section 4 discusses extensions of the presented setting to ambiguity sets with infinitely many probability measures and to continuous state spaces. In Section 5 we provide numerical experiments showcasing the tractability of our $Q$-learning algorithm. The proofs are reported in Section 6.
@@ 2. Setting and specification of the problem
In this section we introduce the underlying framework which will be employed to establish a robust Q-learning algorithm for finite ambiguity sets.
@@ 2.1. Setting
where for all $k\in\{1,\dots,N\}$ and for all $(x,a)\in\mathcal{X}\times A$ we have that $\mathbb{P}^{(k)}(x,a)$ is a probability measure, i.e., $\mathbb{P}^{(k)}(x,a)\in\mathcal{M}_{1}(\mathcal{X})$. Note that in the degenerate case where $\mathcal{P}(x,a)$ contains only a single probability distribution, we are facing a classical Markov decision process (i.e. a non-robust one), compare, e.g., [8].
@@
where the notation $\mathbb{P}=\delta_{x}\otimes\mathbb{P}_{0}\otimes\mathbb{P}_{1}\otimes...\in\mathfrak{P}_{x,a}$ abbreviates the infinite concatenation of the conditional probability distributions:
@@
The construction of ambiguity sets $\mathfrak{P}_{x,\textbf{a}}$ provided in (2.1) and (2.2) is in the related literature often referred to as a $(s,a)$-rectangular ambiguity set, (compare [18] and [52]) and enables to obtain a dynamic programming principle of the associated Markov decision problem which allows to derive tractable numerical solution methods, compare Section 2.4.
@@ 2.3. Specification of the optimization problem
After execution of an action $a_{t}(X_{t})$, the agent receives a feedback on the quality of the chosen action in terms of a reward $r(X_{t},a_{t}(X_{t}),X_{t+1})$ where $r$ denotes some reward function $r:\mathcal{X}\times A\times\mathcal{X}\rightarrow\mathbb{R}$. The robust optimization problem consists then, for every initial state $x\in\mathcal{X}$, in maximizing the expected value of $\sum_{t=0}^{\infty}\alpha^{t}r(X_{t},a_{t}(X_{t}),X_{t+1})$ under the worst case measure from $\mathfrak{P}_{x,\textbf{a}}$ over all possible policies $\textbf{a}\in\mathcal{A}$, where $\alpha\in\mathbb{R}$ is some discount factor accounting for time preferences of rewards. Therefore, the value function
@@ 2.4. Dynamic Programming
To solve (2.3) directly is typically a non-feasible optimization problem as it means to find directly (infinite-dimensional) solutions over the whole time horizon. However, due to its time-homogeneous structure, solving the infinite time horizon problem can be reduced to a one time step problem which turns out to be tractable. To see this, we consider the one step optimization problem
@@
(2.5) $\mathcal{X}\times A\ni(x,a)\mapsto Q^{*}(x,a):=\min_{\mathbb{P}\in\mathcal{P}(x,a)}\mathbb{E}_{\mathbb{P}}\left[r(x,a,X_{1})+\alpha V(X_{1})\right],$
allowing us to interpret $Q^{*}(x,a)$ as the quality of executing action $a$ when in state $x$ (under the worst case measure) with the best possible action leading to $\mathcal{T}V(x)$. By using the above definitions we then obtain the dynamic programming equation $\mathcal{T}V=V$ explaining the notion of quality. This fixed-point equation is also famously known in its non-robust formulation as Bellman equation, compare e.g. [13, p.164] for a non-robust version.
@@ 3. Robust Q-Learning Algorithm
4. Extensions
We acknowledge that the assumption on a finite state space and finite ambiguity sets imposed in Section 2 might be too restrictive for most real-world applications.
To address this drawback, in this section, we indicate by studying two extensions of the presented approach its flexibility and showcase how to apply our $Q$-learning framework in more general settings relevant for practical applications in less contrived settings.
For non-robust $Q$-learning ([51]), the obstacle of finite state spaces has been overcome mainly through the use of function approximation techniques, see, e.g. [5], [26], [27] and [45]. In Section 4.2 we outline a similar extension for our setting.
Moreover, in Section 4.1, we discuss how infinitely large ambiguity sets can be approximated by finite ambiguity sets to which the setting of Section 2 applies.
4.1. Approximation of infinite ambiguity sets
Consider some possibly infinitely large ambiguity set of probability measures
[…]
for which we want to compute its $Q$-value function $Q^{*}(x,a)=\inf_{\mathbb{P}\in\mathcal{P}(x,a)}\mathbb{E}_{\mathbb{P}}\left[r(x,a,X_{1})+\alpha V(X_{1})\right]$. Moreover, assume that $\mathcal{P}(x,a)$ fulfils [31, Assumption 2.2] so that minimizers of the corresponding value function exist (compare [31, Theorem 2.7]). Under these assumptions, we can derive the following Lemma 4.1 that allows to approximate the corresponding $Q$-value function by the $Q$-value function of an approximating sequence of finite ambiguity sets as defined in (2.1), justifying in return to apply Algorithm 1 to compute an approximation of the $Q$-value function associated to $\mathcal{P}(x,a)$.
Lemma 4.1.
If there exists a sequence of finite ambiguity sets $\left(\mathcal{P}^{(n)}(x,a)\right)_{n\in\mathbb{N}}\subset\mathcal{P}(x,a)$ such that for all $(x,a)\in\mathcal{X}\times A$ and all $\mathbb{P}\in\mathcal{P}(x,a)$ there exists a sequence $(\mathbb{P}^{(n)})_{n\in\mathbb{N}}$ such that [^2] Note that the conditions in (4.1) are equivalent to convergence in the Wasserstein-$1$-distance, compare, e.g. [47, Definition 6.8 (i)].
(4.1) $\mathbb{P}^{(n)}\rightarrow\mathbb{P}\text{ weakly, and }\int_{\mathcal{X}}|x|{\mathop{}\!\mathrm{d}}\mathbb{P}^{(n)}(x)\rightarrow\int_{\mathcal{X}}|x|{\mathop{}\!\mathrm{d}}\mathbb{P}(x)\text{ as }n\rightarrow\infty,$
then we also have
$\lim_{n\rightarrow\infty}Q^{(n)}(x,a)$ $:=\lim_{n\rightarrow\infty}\inf_{\mathbb{P}\in\mathcal{P}^{(n)}(x,a)}\mathbb{E}_{\mathbb{P}}\left[r(x,a,X_{1})+\alpha V^{(n)}(X_{1})\right]$
$=\inf_{\mathbb{P}\in\mathcal{P}(x,a)}\mathbb{E}_{\mathbb{P}}\left[r(x,a,X_{1})+\alpha V(X_{1})\right]=:Q(x,a).$
where $V^{(n)}$ denotes the value function as defined in (2.3) associated to the finite ambiguity set $\mathcal{P}^{(n)}(x,a)$.
Example 4.2.
Consider the state space $\mathcal{X}=\{1,\dots,N\}$ for some $N\in\mathbb{N}$. Then, for any two functions $\underline{p}:\mathcal{X}\rightarrow\mathbb{R}_{+},\overline{p}:\mathcal{X}\rightarrow\mathbb{R}_{+}$ with $\underline{p}(\cdot)\leq\overline{p}(\cdot)$, the ambiguity set of binomial distributions
$\mathcal{X}\times A\ni(x,a)\twoheadrightarrow\mathcal{P}(x,a):=\{\operatorname{Bin}(N,p),p\in[\underline{p}(x),~{}\overline{p}(x)]\}$
fulfils [31, Assumption 2.2] and can be approximated in the sense of Lemma 4.1 by
$\mathcal{P}^{(n)}(x,a):=\left\{\operatorname{Bin}(N,p),p\in\{p_{0},\dots,p_{n}\}\text{ with }p_{i}=\underline{p}(x)+\frac{i}{n}\left(\overline{p}(x)-\underline{p}(x)\right),~{}~{}i=0,\dots,n\right\}.$
The proof is reported in the appendix. This means, by Lemma 4.1, to approximate the $Q$-value function associated to ambiguity set $\mathcal{P}(x,a)$, we can compute the $Q$-value function associated to ambiguity set $\mathcal{P}^{(n)}(x,a)$ using Algorithm 1 for a sufficiently large $n$.
4.2. Continuous state space
Algorithm 1 assumes a finite state space and can therefore not be applied to problems formulated on a continuous state space. Howeover, if we have a continuous state space $\mathcal{X}\subset\mathbb{R}^{d}$, finite action space $A\subset\mathbb{R}^{m}$, and a finite ambiguity set
then we can pursue a similar approach as proposed in [27] or [45] for the non-robust setting.
To this end, we parameterize the $Q$-value function by some approximating function, where a popular choice int he field of reinforcement learning are neural networks $\mathcal{X}\times\mathcal{A}\ni(x,a)\mapsto Q_{\theta}(x,a)$ parameterized by some $\theta$. The objective is then to minimize the error between $Q_{\theta}(x,a)$ and
$\min_{k=1,\dots,N}\mathbb{E}_{\mathbb{P}^{(k)}(x,a)}\left[r(x,a,X_{1})+\alpha\max_{b\in A}Q_{\theta}(X_{1},b)\right]$
on a batch of state action pairs, where the expectation is approximated via Monte-Carlo simulation.
Hence, we sample a batch of state action pairs $(x^{i},a^{i})$, $i=1,\dots,B$ for batch size $B\in\mathbb{N}$ and minimize the loss function
$\sum_{i=1}^{B}\left(Q_{\theta}(x^{i},a^{i})-\min_{k=1,\dots,N}\frac{1}{N_{\rm MC}}\sum_{j=1}^{N_{\rm MC}}\left[r(x^{i},a^{i},X^{k,j})+\alpha\max_{b\in A}Q_{\theta}(X^{k,j},b)\right]\right)^{2}$
via stochastic gradient descent, where $X^{k,j}\sim\mathbb{P}^{(k)}(x^{i},a^{i})$ for $j=1,\dots,N_{\rm MC}$.
Note that adaptions such as double Q-learning ([45]) are easily possible also in the present robust setting.
5. Numerical Experiments
@@ 5.1. Coin Toss
As in [31, Example 4.1] we consider an agent playing a coin toss game.
@@
At each time step $t$ the agent can make a bet whether the sum of the next throw strictly exceeds the previous sum (i.e., $X_{t+1}>X_{t}$), or whether it is strictly smaller (i.e., $X_{t+1}<X_{t}$). If the agent is correct, she gets $1\$$, however if the agent is wrong she has to pay $1\$$. The agent also has the possibility not to play. We model this by considering the following reward function:
@@
which are Wasserstein-balls around the reference probability $\operatorname*{Bin}(10,0.5)$ with radii $1$ and $2$, respectively. As one can see, e.g., from [29, Equation (4.2)], $\mathcal{P}^{3}$ is the smallest Wasserstein-ball around the reference measure that contains $\operatorname*{Bin}(10,0.6)$, whereas $\mathcal{P}^{4}$ is the smallest Wasserstein-ball containing $\operatorname*{Bin}(10,0.3)$.
@@ 5.2. Stock investing
This example is a variant of [31, Example 4.3] showcasing the use of tailored ambiguity sets which are not possible to construct and use in the same manner with Wasserstein or Entropic ambiguity sets as they are considered in [7], [23], and [48].
@@ 6.1. Auxiliary Results
Let $10<\alpha<$ 1, and let the ambiguity set $\mathcal{P}$ be defined in (2.1). Then the following fixed-point equation holds true for the optimal $Q$-value function defined in (2.5)
@@
This follows directly by definition of $Q^{*}$ and by Proposition 2.1, compare also [29, Proof of Lemma 17]. ∎
@@
The proof follows is analogue to [29, Proof of Lemma 18]. ∎
The following lemma is a result from stochastic approximation theory adjusted to our setting, compare [29, Lemma A.1] for the presented formulation adjusted to the underlying setting. We refer also to [39, Lemma 1], [15] [19, Theorem 1], [41, Lemma 12], and [46, Lemma 3] for further reference.
@@
6.2. Proofs of the results from Section 2, Section 3, and Section 4
@@
Note that the equality $\max_{a\in A}Q^{*}(x,a)=\mathcal{T}V(x)$ directly follows by definition of the operator $\mathcal{T}$, and it only remains to show $\mathcal{T}V=V$. To this end, we aim at applying [31, Theorem 3.1] with $p=0$ and verify its assumptions, i.e., we verify [31, Assumption 2.2] and [31, Assumption 2.4] by showing the following four properties.
@@
To see that (i) holds, we note that $\mathcal{P}$ is by definition non-empty and since the set $\mathcal{P}(x,a)$ is finite it is also compact for all $(x,a)\in\mathcal{X}\times A$. To show the upper hemicontinuity, we consider some $(x,a)\in\mathcal{X}\times A$ and we let $(x_{n},a_{n},\mathbb{P}_{n})_{n\in\mathbb{N}}\subseteq\operatorname{Gr}(\mathcal{P}):=\{(x,a,\mathbb{P})~{}|~{}(x,a)\in\mathcal{X}\times A,\mathbb{P}\in\mathcal{P}(x,a)\}$ be a sequence such that $(x_{n},a_{n})\rightarrow(x,a)\in\mathcal{X}\times A$ as $n\rightarrow\infty$. Since both $\mathcal{X}$ and $A$ are finite, we have $(x_{n},a_{n})=(x,a)$ for $n$ large enough. Hence, $\mathbb{P}_{n}\in\mathcal{P}(x,a)$ for $n$ large enough. As $\mathcal{P}(x,a)$ is finite we also have that $\mathbb{P}_{n}=\mathbb{P}^{(i)}(x,a)$ for infinitely many indices $n$ for some $i\in\{1,\dots,N\}$. Hence, we can find a subsequence $(\mathbb{P}_{n_{k}})_{k\in\mathbb{N}}$ such that $\mathbb{P}_{n_{k}}\rightarrow\mathbb{P}^{(i)}(x,a)\in\mathcal{P}(x,a)$ weakly as $k\rightarrow\infty$ and the upper hemicontinuity of $\mathcal{P}$ follows by the characterization provided in [2, Theorem 17.20].
To show the lower hemicontinuity let $(x,a)\in\mathcal{X}\times A$, $\mathbb{P}\in\mathcal{P}(x,a)$ and consider a sequence $(x_{n},a_{n})_{n\in\mathbb{N}}\subseteq\mathcal{X}\times A$ with $(x_{n},a_{n})\rightarrow(x,a)$ as $n\rightarrow\infty$. By definition of $\mathcal{P}$, there exists some $i_{\mathbb{P}}\in\{1,\dots,N\}$ such that $\mathbb{P}^{(i_{\mathbb{P}})}(x,a)=\mathbb{P}$. We use this observation to define
$\mathbb{P}_{n}:=\mathbb{P}^{(i_{\mathbb{P}})}(x_{n},a_{n})\in\mathcal{P}(x_{n},a_{n}),~{}n\in\mathbb{N},$
@@
$k_{t}^{*}\in\operatorname*{arg\,min}_{k=1,...,N}\mathbb{E}_{\mathbb{P}^{(k)}(X_{t},a_{t}(X_{t}))}\left[r(X_{t},a_{t}(X_{t}),X_{t+1})+\alpha\max_{b\in A}Q_{t}(X_{t+1},b)\right].$
@@
Proof of Lemma 4.1.
Let $(x,a)\in\mathcal{X}\times A$, then, by [31, Theorem 2.7 (i)], there exists some minimizing measure $\mathbb{P}^{*}\in\mathcal{P}(x,a)$ such that
$\inf_{\mathbb{P}\in\mathcal{P}(x,a)}\mathbb{E}_{\mathbb{P}}\left[r(x,a,X_{1})+\alpha V(X_{1})\right]=\mathbb{E}_{\mathbb{P}^{*}}\left[r(x,a,X_{1})+\alpha V(X_{1})\right].$
Take some sequence $(\mathbb{P}^{(n)})_{n\in\mathbb{N}}$ with $\mathbb{P}^{(n)}\in\mathcal{P}^{(n)}(x,a)$ for all $n\in\mathbb{N}$ such that $\mathbb{P}^{(n)}\rightarrow\mathbb{P}^{*}$ weakly as $n\rightarrow\infty$ and such that $\int_{\mathcal{X}}|x|{\mathop{}\!\mathrm{d}}\mathbb{P}^{(n)}(x)\rightarrow\int_{\mathcal{X}}|x|{\mathop{}\!\mathrm{d}}\mathbb{P}(x)$. Then, since $\mathcal{X}$ is finite, we directly obtain
$\inf_{\mathbb{P}\in\mathcal{P}(x,a)}\mathbb{E}_{\mathbb{P}}\left[r(x,a,X_{1})+\alpha V(X_{1})\right]$ $=\lim_{n\rightarrow\infty}\mathbb{E}_{\mathbb{P}^{(n)}}\left[r(x,a,X_{1})+\alpha V(X_{1})\right]$
$\geq\lim_{n\rightarrow\infty}\inf_{\mathbb{P}\in\mathcal{P}^{(n)}(x,a)}\mathbb{E}_{\mathbb{P}}\left[r(x,a,X_{1})+\alpha V(X_{1})\right].$
But note that, since $\mathcal{P}^{(n)}(x,a)\subset\mathcal{P}(x,a)$, we also have
$\inf_{\mathbb{P}\in\mathcal{P}(x,a)}\mathbb{E}_{\mathbb{P}}\left[r(x,a,X_{1})+\alpha V(X_{1})\right]\leq\inf_{\mathbb{P}\in\mathcal{P}^{(n)}(x,a)}\mathbb{E}_{\mathbb{P}}\left[r(x,a,X_{1})+\alpha V(X_{1})\right]$
showing that
(6.9) $\inf_{\mathbb{P}\in\mathcal{P}(x,a)}\mathbb{E}_{\mathbb{P}}\left[r(x,a,X_{1})+\alpha V(X_{1})\right]$ $=\lim_{n\rightarrow\infty}\inf_{\mathbb{P}\in\mathcal{P}^{(n)}(x,a)}\mathbb{E}_{\mathbb{P}}\left[r(x,a,X_{1})+\alpha V(X_{1})\right]$
$=\lim_{n\rightarrow\infty}\mathbb{E}_{\mathbb{P}^{(n)}}\left[r(x,a,X_{1})+\alpha V(X_{1})\right].$
Let $V_{\varepsilon_{n}}(x)$ denote the value function defined as in (2.3) associated to the ambiguity set $\mathcal{P}_{\varepsilon_{n}}(x,a):=\{\mathbb{P}\in\mathcal{M}_{1}(\mathcal{X})~{}|~{}d_{W}(\mathbb{P},\mathbb{P}^{*})\leq\varepsilon_{n}\}$ where $\varepsilon_{n}=d_{W}(\mathbb{P}^{(n)},\mathbb{P}^{*})$, for $d_{W}$ denoting the $1$-Wasserstein-distance defined by
$d_{W}(\mathbb{P}^{(n)},\mathbb{P}^{*}):=\inf_{\pi\in\Pi(\mathbb{P}^{(n)},\mathbb{P}^{*})}\int_{\mathcal{X}\times\mathcal{X}}\|x-y\|{\mathop{}\!\mathrm{d}}\pi(x,y),$
[…]
(6.10) $|V^{*}(x)-V^{(n)}(x)|\leq|V^{*}(x)-V_{\varepsilon_{n}}(x)|+|V_{\varepsilon_{n}}(x)-V^{(n)}(x)|\leq C\varepsilon_{n}$
for some constant $C$ independent of $x\in\mathcal{X}$, since both $\mathbb{P}^{*}$ and $\mathbb{P}^{(n)}$ are by definition inside a Wasserstein-ball centered at $\mathbb{P}^{*}$ with radius $\varepsilon_{n}$. Hence, we get together with (6.9)
$\lim_{n\rightarrow\infty}\inf_{\mathbb{P}\in\mathcal{P}^{(n)}(x,a)}\mathbb{E}_{\mathbb{P}}\left[r(x,a,X_{1})+\alpha V^{(n)}(X_{1})\right]$
$=\lim_{n\rightarrow\infty}\mathbb{E}_{\mathbb{P}^{(n)}}\left[r(x,a,X_{1})+\alpha V(X_{1})\right]+\mathbb{E}_{\mathbb{P}^{(n)}}\left[\alpha\left(V^{(n)}(X_{1})-V(X_{1})\right)\right]$
$=\inf_{\mathbb{P}\in\mathcal{P}(x,a)}\mathbb{E}_{\mathbb{P}}\left[r(x,a,X_{1})+\alpha V(X_{1})\right]$
by (6.10) since $\varepsilon_{n}=d_{W}(\mathbb{P}^{(n)},\mathbb{P}^{*})\rightarrow 0$ as $n\rightarrow\infty$ which follows from [47, Definition 6.8 (i)] and by the assumption that $\mathbb{P}^{(n)}\rightarrow\mathbb{P}$ weakly as well as $\int_{\mathcal{X}}|x|{\mathop{}\!\mathrm{d}}\mathbb{P}^{(n)}(x)\rightarrow\int_{\mathcal{X}}|x|{\mathop{}\!\mathrm{d}}\mathbb{P}(x)$. ∎
Proof of Example 4.2.
We first show that [31, Assumption 2.2] is fulfilled for the weak topology (i.e., $p=0$ in the notation of [31]). This means we want to show that
$\mathcal{X}\times A\ni(x,a)\twoheadrightarrow\mathcal{P}(x,a):=\{\operatorname{Bin}(N,p),p\in[\underline{p}(x),\overline{p}(x)]\}$
is non-empty, compact-valued, and continuous.
The non-emptiness follows from requiring $\underline{p}(\cdot)\leq\overline{p}(\cdot)$.
For the compactness let $(\mathbb{P}^{(n)})_{n\in\mathbb{N}}\subseteq\mathcal{P}(x,a)$, then we have for all $n\in\mathbb{N}$ the representation $\mathbb{P}^{(n)}=\operatorname{Bin}(N,p^{(n)})$ for some $p^{(n)}\in[\underline{p}(x),\overline{p}(x)]$. Hence, by the Bolzano–Weierstrass theorem there exists a convergent subsequence $p^{(n_{k})}\rightarrow p\in[\underline{p}(x),\overline{p}(x)]$ as $k\rightarrow\infty$ which also implies $\mathbb{P}^{(n_{k})}\rightarrow\mathbb{P}=\operatorname{Bin}(N,p)$ weakly.
For the continuity we use the characterization of upper and lower hemicontinuity provided in [2, Theorem 17.20 and Theorem 17.21]. For the upper hemicontinuity take a sequence $(x^{(n)},a^{(n)})\subseteq\mathcal{X}\times A$ with $(x^{(n)},a^{(n)})\rightarrow(x,a)\in\mathcal{X}\times A$ as $n\rightarrow\infty$, and consider a sequence $(\mathbb{P}^{(n)})_{n\in\mathbb{N}}$ with $\mathbb{P}^{(n)}\in\mathcal{P}(x^{(n)},a^{(n)})$ for all $n\in\mathbb{N}$. Then, we can represent $\mathbb{P}^{(n)}=\operatorname{Bin}(N,p^{(n)})$ for some $p^{(n)}\in[\underline{p}(x^{(n)}),\overline{p}(x^{(n)})]$ and for $n$ large enough $x^{(n)}=x$, hence $p^{(n)}\in[\underline{p}(x),\overline{p}(x)]$, i.e., $\mathbb{P}^{(n)}\in\mathcal{P}(x,a)$ showing the upper hemicontinuity. To show the lower hemicontinuity, take a sequence $(x^{(n)},a^{(n)})\subseteq\mathcal{X}\times A$ with $(x^{(n)},a^{(n)})\rightarrow(x,a)\in\mathcal{X}\times A$ and let $\mathbb{P}\in\mathcal{P}(x,a)$, then $\mathbb{P}=\operatorname{Bin}(N,p)$ for some $p=\lambda\underline{p}(x)+(1-\lambda)\overline{p}(x)$ with $\lambda\in[0,1]$. Define $\mathbb{P}^{(n)}=\operatorname{Bin}(N,p^{(n)})$ for $p^{(n)}=\lambda\underline{p}(x^{(n)})+(1-\lambda)\overline{p}(x^{(n)})$ implying $\mathbb{P}^{(n)}\in\mathcal{P}(x^{(n)},a^{(n)})$. Then, by definition $p^{(n)}\rightarrow p$ and hence $\operatorname{Bin}(N,p^{(n)})\rightarrow\operatorname{Bin}(N,p)$ as $n\rightarrow\infty$ weakly, showing the lower hemicontinuity.
Next, we show that $\mathcal{P}^{(n)}(x,a)$ fulfils the assumptions of Lemma 4.1. To this end, let $(x,a)\in\mathcal{X}\times A$ and $\mathbb{P}\in\mathcal{P}(x,a)$, i.e., $\mathbb{P}=\operatorname{Bin}(N,p)$ for some $p=\lambda\underline{p}(x)+(1-\lambda)\overline{p}(x)$ with $\lambda\in[0,1]$. Define $\mathbb{P}^{(n)}\in\mathcal{P}^{(n)}(x,a)$ by $\mathbb{P}^{(n)}:=\operatorname{Bin}(N,p_{i})$ where $i\in\{1,\dots,n\}$ such that $|\frac{i}{n}-\lambda|$ is minimal, implying that $p_{i}\rightarrow p$ as $n\rightarrow\infty$. Then, we have for all continuous and bounded functions $f:\mathcal{X}\rightarrow\mathbb{R}$ that
$\int_{\mathcal{X}}f(x){\mathop{}\!\mathrm{d}}\mathbb{P}^{(n)}(x)$ $=\sum_{k=0}^{N}{N\choose k}p_{i}^{k}(1-p_{i})^{k}f(k)\rightarrow\sum_{k=0}^{N}{N\choose k}p^{k}(1-p)^{k}f(k)=\int_{\mathcal{X}}f(x){\mathop{}\!\mathrm{d}}\mathbb{P}(x)$
as $n\rightarrow\infty$, i.e., $\mathbb{P}^{(n)}\rightarrow\mathbb{P}$ weakly. Analogously, we obtain $\int_{\mathcal{X}}|x|{\mathop{}\!\mathrm{d}}\mathbb{P}^{(n)}(x)\rightarrow\int_{\mathcal{X}}|x|{\mathop{}\!\mathrm{d}}\mathbb{P}(x)$. ∎
Acknowledgements
J.S. gratefully acknowledges financial support by the NUS Start-Up Grant Tackling model uncertainty in Finance with machine learning.